# Non-perturbative approach for scalar particle production in Higgs-$R^{2}$ inflation

### v1

@@ II.1 Background dynamics
$\dot{H}$ $=$ −$-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (20)
One has to numerically solve the system of equations (17), (18), (19) and (20) in order to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given a suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (14) inside one of the valleys of the potential. In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys is an effective single-field description. Then the scalaron field should fulfill the single-field equation
@@
which is intregrable by a quadrature together with the Friedman equations:
@@
is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation. It is clear from slow-roll solutions that during inflation we have the usual quasi De Sitter solutions for $\phi$, $a$ and $H$. The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
@@
III Free scalar fields in the Einstein frame
In this section we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initial assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflatonary sector. In the Jordan frame our model is giving by
$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,,$ (29)
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. The action of such field is giving by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (30)
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to gravity sector and may take any value; however, studies based on electroweak vacuum stability [41] suggest that the value of non-minimal coupling constant should not exceed the value $10$. Moreover, there are two interest special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called minimal coupling case and the value $\xi_{\chi}=1/6$ is called conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if in addition $m_{\chi}=0$. Although there are no direct coupling between matter fields and the inflatonary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field is present during inflation and after inflation, and may occur gravitational particle production process. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also ocurr as long as the matter fields are not conformally invariant, namely, if the scalar fied is massive $m_{\chi}\neq 0$. These kind of models which in the scalar fields interacts only gravitationally is specially interesting since it can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications in the reheating process [17, 42, 43] and gravitational dark matter production [44, 45, 46, 12, 11, 9, 47].
@@ III Free scalar fields in the Einstein frame
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involve the scalar field $\tilde{\chi}$ leading to another Einstein frame in order to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is an spectator field and both scalaron field $\phi$ and Higgs field are responsible of the dynamics during and after inflation. One can see that the conformal transformation (5) induces interaction terms between the scalaron $\phi$, the Higgs field $h$ and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
@@
$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi\right.\\[8.5359pt] \left.+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi\right]\,.$ (32)
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$). One can see this fact from the decay ratio of the scalaron into a pair of light scalar particles during reheating epoch [8], [48]
$\Gamma_{\phi\to\chi\chi}=\left(1-6\xi_{\chi}+\dfrac{2m_{\chi}^{2}}{M^{2}}\right)^{2}\,\dfrac{M^{3}\sqrt{1-4m_{\chi}^{2}/M^{2}}}{192\pi M_{p}^{2}}\,.$
@@
If $\xi_{\chi}=1/6$ and $m_{\chi}\to 0$, the decay ratio vanishes. In addition, if the mass of scalar field $\chi$ is $m_{\chi}>M/2$ where $M$ is scalaron mass term, the scalaron decay into a pair of $\chi$ particles is forbidden. Then, the perturbative production of $\chi$ particles is allowed only in the case of light scalars with mass $m_{\chi}<M/2$. However, it can also occur a scalar particle production by non perturbative effects by time-dependent background dynamics of $\phi$, $h$ and $a$ for a spectator field with mass $m_{\chi}>M/2$. In order to study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation. Varying the action (31) respect to $\tilde{\chi}$ yields the following equation of motion
(□−$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}\\[5.69046pt] +\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (34)
@@
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, specially after inflation when both fields $\phi$ and $h$ exhibit a strongly interacting anharmonic oscillatory behavior, as one can see in the figure 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (21).
IV Mode evolution
@@
In this section, we numerically solve the equation of motion (34) along with the background equations (17), (18), (I) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$. The equation of motion for $\tilde{\chi}$ considering a FLRW universe (I) can be written in terms of the cosmic time $t$ as
@@
Particle production in the early universe is a process of quantum field theory in curved spacetime, where the gravitational field is not quantized, but the free scalar field $\chi$ is quantized. Therefore, the production of $\chi$ particles is a quantum effect [3, 4, 5]. We proceed by quantizing the spectator field in the Heisenberg picture so that the $\chi$ field and its conjugate momentum $\pi_{\chi}=\dot{\chi}$ are treated as field operators. In Fourier space the rescaled field can be written as
@@
$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (42)
The equation for $X_{k}$ describes an oscillator with time-dependent frequency $\omega_{k}^{2}(t)$, hence to track the evolution of the modes $X_{k}$ during and after inflation, it is necessary to analyze $\omega_{k}^{2}(t)$. One can see that during the quasi De Sitter phase during inflation when the effective single field inflation is valid, the effective mass is simplified because of $\dot{H}\to 0$. In this case one can see that the effective mass takes the form
$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}\\[5.69046pt] +3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (43)
We can see that the mass term is exponentially suppressed during inflation when $\phi\gg M_{p}$, making the mass of the spectator field much smaller than the Hubble parameter at the end of inflation, $m_{\chi}\ll H_{\mathrm{end}}$. However, after inflation, when the scalaron takes values $\phi\ll M_{p}$, the free scalar $\tilde{\chi}$ becomes superheavy, $m_{\chi}\gg H_{\mathrm{end}}$, due to the exponential dependence on the scalaron $\phi$. Then at the beginning of inflation, when $t\ll t_{\mathrm{end}}$ the Ricci scalar is $R\simeq 12H^{2}$ and the dominant terms in (41) are $k^{2}/a^{2}$, $H^{2}$ and $12\xi_{\chi}H^{2}$. However, towards the end of inflation, the term $\dot{H}$ cannot be neglected, as after inflation, $\dot{H}$ and the background oscillations dominate the dynamics. The term $V_{\phi}$ regulates the oscillations of the effective mass after inflation, but it is small compared to $H^{2}$ and $\dot{H}$. The contribution of each term depends on the moment at which it is considered. The mass term benefits from exponential growth when $\phi\ll M_{p}$, as we anticipated earlier. Therefore, after inflation, the dominant terms are the mass term and $\dot{H}$. Additionally, if the coupling constant is $\xi_{\chi}<1/6$, the effective mass could be negative during inflation. In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to a exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability. This regime may result in a very efficient amplification of particle production [42, 33].
Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}M$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing. We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
However, for large enough modes $k$, namely
@@
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating). One can also observe growth due to the exponential dependence of the scalaron in the mass term toward the end of inflation.
To numerically solve equation (40) along with the background equations, we must specify suitable intial conditions that ensure the modes are within the horizon at early times. In the asymptotic limit $t\to 0$ and according to the behavior of $\omega_{k}^{2}(t)$ at the beginning of inflation, the dominant term is $k/a$. In this limit, the modes are sub-horizon $k\gg aH$, and the modes exhibit Minkowski-like behavior. This motivates choosing the Bunch-Davies vacuum as the initial condition [14]
@@
V Description of particle production
Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wave number $k$, it is turn to analyze the non-perturbative particle production. The production of $\chi$ particles is due to the time dependence of the background variables, in particular the non-adiabatic behaviour of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3]. The adiabaticity condition is given by the adiabatic paramater $A_{k}$ defined as
Figure 4: An example of non-adiabatic mode evolution leading the gravitational particle production of $\chi$ particles described by the adiabaticity parameter $|A_{k}(t)|$ for $m_{\chi}=10^{-2}M$ and $\xi_{\chi}=1/6$. At the end of inflation, one see an abruptly increase on $|A_{k}(t)|$ and non-adiabaticity is induced, leading gravitational particle production $n_{k}(t)\neq 0$. At late times, when the stabilization occurs, we recover the adiabaticity evolution.

### v2

@@ 2.1 Background dynamics
$\dot{H}$ $=$ $-\dfrac{1}{2M_{p}^{2}}\,\left(\dot{\phi}^{2}+e^{-\alpha\phi}\dot{h}^{2}\right)\,.$ (19)
One has to numerically solve the system of equations (16), (17), (18) and (19) to explore the inflationary dynamics of the model in the Einstein frame given some combination values of the parameters $\xi_{h}\,,\xi_{s}\,,\lambda$ and given suitable initial conditions for $\phi(t_{0})$, $h(t_{0})$. The inflation occurs when $\phi\gg M_{p}$ and the value of the auxiliary field $h$ is given by (13) inside one of the valleys of the potential. In the figure 2 we present the oscillatory regime of both fields after inflation in terms of cosmic time for the initial conditions $\phi_{0}=5.7M_{p}$ and $h_{0}\approx 0.096M_{p}$. However, according to the argument we have presented above, during inflation the isocurvature effects are absent and the dynamics inside the valleys are an effective single-field description. Then the scalaron field should fulfill the single-field equation
@@
which can be integrated by quadrature together with the Friedman equations
@@
is the time that the effective inflation ends and $H_{0}$ is the scale Hubble parameter during inflation, with $\phi_{0}\sim 5.7M_{p}$ is the value of $\phi$ at the beginning of inflation. It is clear from slow-roll solutions that during inflation we have the usual quasi-De Sitter solutions for $\phi$, $a$, and $H$. The effective single-field inflation ends when the slow roll parameter $\epsilon$ is $\epsilon(t_{\mathrm{end}})=1$, which happens at the final values
@@
3 Scalar fields in the Einstein frame
In this section, we study spectator fields present in the Higgs-$R^{2}$ model such as free scalar fields, initially assuming no direct coupling between the matter fields and the inflationary sector in the Jordan frame. In other words, we get started in the Jordan frame and after conformal transformation, the matter sector is gravitationally coupled to the inflationary sector. In the Jordan frame our model is given by
$S_{J}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}H^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]+S_{\mathrm{matt}}[\psi_{i}]\,$
where $\psi_{i}$ are the different matter fields that are present during and after inflation. In this work, we will consider a massive spectator scalar field $\chi$ that only interacts gravitationally. The action of such a field is given by
$S_{J}[\chi]=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left(-\dfrac{1}{2}g^{\mu\nu}_{J}\partial_{\mu}\chi\partial_{\nu}\chi-\dfrac{m_{\chi}^{2}}{2}\chi^{2}\dfrac{1}{2}\,\xi_{\chi}\,R_{J}\,\chi^{2}\right)\,.$ (29)
The dimensionless parameter $\xi_{\chi}$ represents the non-minimal coupling to the gravity sector and may take any value; however, studies based on electroweak vacuum stability [40] suggest that the value of non-minimal coupling constant should not exceed $10$. Moreover there are two interesting special cases: $\xi_{\chi}=0$ and $\xi_{\chi}=1/6$. The value $\xi_{\chi}=0$ is called the case of minimal coupling. while $\xi_{\chi}=1/6$ corresponds to conformal coupling. The first case corresponds to a free scalar field theory and the second one ensures the conformal invariance of the model if, in addition, $m_{\chi}=0$.
Although there is no direct coupling between matter fields and the inflationary sector in the first case, both scalaron $\phi$ and Higgs field $h$ gravitationally interact with every matter field present during inflation and after inflation, and a gravitational particle production process may occur. In the second one, free scalar $\chi$ is coupled to the Ricci scalar and the particle production may also occur as long as the matter fields are not conformally invariant, namely, if the scalar field is massive $m_{\chi}\neq 0$. These kinds of models, in which the scalar fields interact only gravitationally, are especially interesting. They can explain the particle production process, with the gravitational field responsible for creating $\chi$ particles. This process of particle creation due to gravitational interactions has implications for the reheating process [17, 41, 42]. It also plays a role in gravitational dark matter production [43, 44, 45, 12, 11, 9, 46].
@@ 3 Scalar fields in the Einstein frame
where $\chi\to\tilde{\chi}=\Psi^{-1}\chi$ is the re-scaled scalar field and the Ricci scalar in the Einstein frame is given by $R=6(\dot{H}+2H^{2})$. We can perform another conformal transformation which involves the scalar field $\tilde{\chi}$ leading to another Einstein frame to remove the non-minimal coupling constant $\xi_{\chi}$. However, we are only interested in the situation where $\chi$ is a spectator field and both scalaron field $\phi$ and Higgs field are responsible for the dynamics during and after inflation. One can see that the conformal transformation (4) induces interaction terms between the scalaron $\phi$, the Higgs field $h$, and the free scalar $\tilde{\chi}$. We can identify the interaction terms from the following action
@@
$S_{\mathrm{int}}=3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\tilde{\chi}\,g^{\mu\nu}\partial_{\mu}\tilde{\chi}\partial_{\nu}\phi+\dfrac{\alpha}{4}\tilde{\chi}^{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right]\,.$ (31)
@@
This action describes the interaction between the background fields $\phi$ and $h$ with the massive spectator field $\tilde{\chi}$. We can see that even with $\xi_{\chi}=0$ this interaction does not disappear, unless the field $\tilde{\chi}$ is conformal invariant ($m_{\chi}=0\,,\xi_{\chi}=1/6$). In the paper [8], the authors study the gravitational production of dark matter in the $R^{2}$ inflation model and show that the scalaron $\phi$ can decay perturbatively in dark matter particles $\chi$ with masses $m_{\chi}<M/2$, where $M\simeq 1.3\times 10^{-5}M_{p}$ is the mass term of the scalaron. For heavy particles, perturbative particle production is kinematically forbidden. However, the production of scalar particles is not limited to perturbative methods; there may be particle production due to the expansion of the universe and the dynamics of the fields involved during the inflationary epoch. In our case, the classical time-dependent background is defined by the evolution of the scale factor $a(t)$ and the evolution of the scalar fields $\phi$ and $h$. That is, the matter-free scalar field $\chi$ propagates in a time-dependent classical background, whose dynamics are given by the evolution of the $\phi$ and $h$ fields of the Higgs-$R^{2}$ model. To study this process it is necessary to find the equation of motion for the spectator field $\tilde{\chi}$ and track the mode evolution during and after inflation. Varying the action (30) respect to $\tilde{\chi}$ yields the following equation of motion
$(\square-e^{-\alpha\phi}\,m_{\chi}^{2}-\xi_{\chi}\,R)\tilde{\chi}+\left[\dfrac{\alpha}{2}\,\left(1-6\xi_{\chi}\right)\square\phi-\dfrac{\alpha^{2}}{4}\left(1-6\xi_{\chi}\right)(\partial\phi)^{2}\right]\tilde{\chi}=0\,.$ (32)
@@
Notice that the background dynamics is explicit in the equation of the spectator field $\tilde{\chi}$ due to the conformal transformation. This induces a time-dependent effective mass term, $m_{\mathrm{eff}}^{2}(t),$ for the spectator field $\tilde{\chi}$. Consequently, in the Einstein frame, the model presents an interaction term with the dynamics of the classical background through the Ricci scalar, the scale factor $a$, and the dynamics of the fields $\phi$ and $h$. As a result, studying the solutions is generally highly complex from an analytical perspective; hence a numerical treatment of the background dynamics is necessary for a comprehensive analysis of mode evolution, especially after inflation when both fields $\phi$ and $h$ exhibit a strong interacting anharmonic oscillatory behavior, as one can see in the figure 2. By numerically determining the background evolution of the scalaron, the Higgs field, and the scale factor, we can compute the Ricci scalar using $R=6(\dot{H}+2H^{2})$, ensuring that the dynamics are well-determined. However, during effective inflation along one of the potential’s valleys, we can approximate the background dynamics using the single-field equation of motion (20).
4 Mode evolution
@@
In this section, we numerically solve the equation of motion (32) along with the background equations (16), (17), (1) to obtain the mode functions $\chi$ for calculating the number density of particles produced, $n_{k}$. The equation of motion for $\tilde{\chi}$ considering a FLRW universe (1) can be written in terms of the cosmic time $t$ as
@@
Particle production in the early universe is a process of quantum field theory in curved spacetime, where the gravitational field is not quantized, but the free scalar field $\chi$ is quantized. Therefore, the production of $\chi$ particles is a quantum effect [3, 4, 5]. We proceed by quantizing the spectator field in the Heisenberg picture so that the $\chi$ field and its conjugate momentum $\pi_{\chi}=\dot{\chi}$ are treated as field operators. In Fourier space, the rescaled field can be written as
@@
$m_{\mathrm{eff}}^{2}(t)=e^{-\alpha\phi}m_{\chi}^{2}+\xi_{\chi}\,R-\dfrac{9}{4}H^{2}-\dfrac{3}{2}\dot{H}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)(V_{\phi}-\alpha M_{p}^{2}\dot{H})\,.$ (40)
The equation for $X_{k}$ describes an oscillator with time-dependent frequency $\omega_{k}^{2}(t)$, hence to track the evolution of the modes $X_{k}$ during and after inflation, it is necessary to analyze $\omega_{k}^{2}(t)$. One can see that during the quasi De Sitter phase during inflation when the effective single field inflation is valid, the effective mass is simplified because of $\dot{H}\to 0$. In this case, one can see that the effective mass takes the form
$m_{\mathrm{eff}}^{2}(t\ll t_{\mathrm{end}})\simeq e^{-\alpha\phi}m_{\chi}^{2}+12\left(\xi_{\chi}-\dfrac{3}{16}\right)H^{2}+3\alpha\left(\xi_{\chi}-\dfrac{1}{6}\right)V_{\phi}\,.$ (41)
We can see that the mass term is exponentially suppressed during inflation when $\phi\gg M_{p}$, making the mass of the spectator field much smaller than the Hubble parameter at the end of inflation, $m_{\chi}\ll H_{\mathrm{end}}$. However, after inflation, when the scalaron takes values $\phi\ll M_{p}$, the free scalar $\tilde{\chi}$ becomes super heavy, $m_{\chi}\gg H_{\mathrm{end}}$, due to the exponential dependence on the scalaron $\phi$. Then at the beginning of inflation, when $t\ll t_{\mathrm{end}}$ the Ricci scalar is $R\simeq 12H^{2}$ and the dominant terms in (39) are $k^{2}/a^{2}$, $H^{2}$ and $12\xi_{\chi}H^{2}$. However, towards the end of inflation, the term $\dot{H}$ cannot be neglected, as after inflation, $\dot{H}$ and the background oscillations dominate the dynamics. The term $V_{\phi}$ regulates the oscillations of the effective mass after inflation, but it is small compared to $H^{2}$ and $\dot{H}$. The contribution of each term depends on the moment at which it is considered. The mass term benefits from exponential growth when $\phi\ll M_{p}$, as we anticipated earlier. Therefore, after inflation, the dominant terms are the mass term and $\dot{H}$. Additionally, if the coupling constant is $\xi_{\chi}<1/6$, the effective mass could be negative during inflation. In such a case, the frequency $\omega_{k}^{2}(t)$ can become negative for super-horizon modes $k\ll aH$, leading to an exponential mode growth $\tilde{\chi}\sim e^{|\omega_{k}|t}$ during inflation due to tachyonic instability. This regime may result in a very efficient amplification of particle production [41, 33]. However, for large enough modes $k$, namely
@@
have positive frequency $\omega_{k}^{2}(t)>0$ even if $m_{\mathrm{eff}}^{2}<0$, which corresponds to sub-horizon regime $k\gg aH$. This occurs when the modes are well inside the horizon at the beginning of inflation. In the figure 3 we show the evolution of $\omega_{k}^{2}(t)$ for a non-minimal coupling $\xi=0$. Before $t_{\mathrm{cross}}$, the frequency is dominated by the decay of the physical mode $k/a\sim e^{-Ht}$. Once the mode crosses the horizon $(k\ll aH)$, $\omega_{k}^{2}(t)$ remains almost constant, and its value is negative since the coefficient of the dominant term $H^{2}$ is negative. For $t\gtrsim t_{\mathrm{end}}$, the frequency increases and oscillates after inflation as the background terms oscillate around the minimum of the potential (preheating).
Figure 3: Time-dependent frequency as a function of cosmic time t for a mass $m_{\chi}=10^{-2}H_{\mathrm{end}}$, a non-minimal coupling $\xi_{\chi}=0$, and $k\gg aH$ for 5 e-folds before horizon crossing. We set $N_{e}=60$ before the end of inflation as the moment when the mode crosses the horizon.
One can also observe growth due to the exponential dependence of the scalaron in the mass term toward the end of inflation. To numerically solve equation (38) along with the background equations, we must specify suitable initial conditions that ensure the modes are within the horizon at early times. In the asymptotic limit $t\to 0$ and according to the behavior of $\omega_{k}^{2}(t)$ at the beginning of inflation, the dominant term is $k/a$. In this limit, the modes are sub-horizon $k\gg aH$, and the modes exhibit Minkowski-like behavior. This motivates choosing the Bunch-Davies vacuum as the initial condition [14]
@@
5 Description of particle production
Once we have established the behavior of mode functions $X_{k}$ during and after inflation for some values of spectator mass, the non-minimal coupling $\xi_{\chi}$ and comoving wavenumber $k$, it turns to analyze the non-perturbative particle production. The production of $\chi$ particles is due to the time dependence of the background variables, in particular, the non-adiabatic behavior of the frequency $\omega_{k}(t)$ that induces a mix among negative- and positive-frequency [3]. The adiabaticity condition is given by the adiabatic parameter $A_{k}$ defined as